# Are AI-Generated Fixes Secure? Analyzing LLM and Agent Patches on SWE-bench

### v2

@@ Abstract
Large Language Models (LLMs) and their agentic frameworks are increasingly adopted to automate software development tasks such as issue resolution and program repair. While prior work has identified security risks in LLM-generated code, most evaluations have focused on synthetic or isolated settings, leaving open questions about the security of these systems in real-world development contexts. In this study, we present the first large-scale security analysis of LLM-generated patches using 20,000+ issues from the SWE-bench dataset. We evaluate patches produced by a standalone LLM (Llama 3.3) and compare them to developer-written patches. We also assess the security of patches generated by three top-performing agentic frameworks (OpenHands, AutoCodeRover, HoneyComb) on a subset of our data. Finally, we analyze a wide range of code, issue, and project-level factors to understand the conditions under which LLMs and agents are most likely to generate insecure code. Our findings reveal that the standalone LLM introduces nearly 9x more new vulnerabilities than developers, with many of these exhibiting unique patterns not found in developers’ code. Agentic workflows also generate a significant number of vulnerabilities, particularly when granting LLMs more autonomy, potentially increasing the likelihood of misinterpreting project context or task requirements. We find that vulnerabilities are more likely to occur in LLM patches associated with a higher number of files, more lines of generated code, and GitHub issues that lack specific code snippets or information about the expected code behavior and steps to reproduce. These results suggest that contextual factors play a critical role in the security of the generated code and point toward the need for proactive risk assessment methods that account for both code and issue-level information to complement existing vulnerability detection tools.
Index Terms:
LLMs, Agentic Frameworks, Security, Issue Resolution
I Introduction
@@
LLMs are rapidly transforming software development, assisting with tasks like code completion, test generation, and bug fixing [1, 2, 3, 4, 5]. They are also increasingly integrated into agentic frameworks that enable autonomous planning and execution of complex tasks, such as issue resolution, where LLM-based agents identify, diagnose, and patch software bugs directly in real-world projects [6, 7, 8]. This widespread adoption of LLMs raises critical concerns about their potential security implications [9, 10, 11, 12]. Studies show that 30-50% of code generated with AI code assistants contains vulnerabilities [13, 14, 15]. As LLMs are increasingly entrusted with making direct commits to production repositories, the risk of introducing vulnerabilities grows. Even a single vulnerable commit, as seen in the Log4Shell exploit, can have far-reaching and costly consequences [16]. Thus, there is an urgent need to understand and mitigate the security risks posed by LLM-based systems.
While several studies have shown that LLMs are prone to generating insecure or low-quality code, most focus on simplified tasks such as coding exercises or identifying vulnerabilities in isolated code snippets [17, 18, 13, 19]. These scenarios lack the complexity and context of issue resolution in real projects, where understanding environment constraints or project-specific dependencies is critical. Additionally, when LLMs are embedded in agentic frameworks that automate repair pipelines, these risks may be amplified [20, 21, 22]. As a result, it remains unclear how secure LLM-generated patches are when fixing real-world issues. Although recent benchmarks like SWE-bench [23] have enabled the assessment of LLMs on real-world software engineering tasks, security-specific analyses in these contexts remain limited. Such evaluations are essential for understanding the circumstances under which vulnerabilities emerge in LLM-generated bug fixes.
In this work, we conduct an empirical study of LLM and agent-generated code in real-world issue resolution tasks, with a focus on security. Using the SWE-bench dataset, we analyze patches generated both by a standalone LLM and three agentic frameworks, comparing them to developer-written patches. Our goal is not only to assess whether LLMs and agentic frameworks introduce security vulnerabilities in realistic settings, but also to understand the conditions under which this occurs. We aim to identify the characteristics of risky LLM and agent-generated contributions and contexts that lead to insecure outputs. We ask the following research questions:
@@
  • RQ1: How secure are the patches produced by standalone LLMs on real-world issue resolution tasks? How does their security compare to developer-written patches?
@@
To answer these questions, we evaluate the vulnerability prevalence, type, and overlaps in: 1) Developer-written patches, 2) Standalone LLM patches (Llama 3.3 70B [24]), and 3) Agentic APR Framework patches (OpenHands [6], AutoCodeRover [8], and Honeycomb [25]). Using a majority voting technique to suppress false positives across three static analysis tools (Semgrep [26], Bandit [27], and CodeQL [28]), we identify the new vulnerabilities introduced through the patches. Lastly, we examine a comprehensive set of of code, issue, and project-level metrics to assess the conditions under which LLMs are most likely to introduce vulnerabilities.
@@
Our findings highlight the security risks of LLM-generated code at scale. When evaluated on 20,000+ GitHub issues, the standalone LLM introduced 185 new vulnerabilities i.e., nearly 9x higher than the 20 new vulnerabilities introduced by developers. On a smaller test set of 2,200+ issues, the agentic frameworks also produced vulnerabilities, with the highest counts observed in the framework that gave the LLM greatest autonomy. Notably, some of the most recurring vulnerabilities introduced by LLMs (CWE-95, CWE-327) were almost entirely distinct from those written by developers (e.g., CWE-732, CWE-377), suggesting that LLMs are not only replicating developers’ errors but also exhibiting new behavioral patterns. We found that vulnerable outputs are more likely when LLM patches involve a higher number of files and generated lines of code. Further, most LLM-generated vulnerabilities were observed in instances involving bug-related issues that do not contain code examples, information about the expected behavior of the code, or steps to reproduce the issue. These findings suggest that the risk of generating vulnerabilities is not only a function of the code produced but also of the context in which the LLM operates. As such, our insights can help inform proactive risk assessments of LLM and agent contributions, complementing the existing vulnerability detection tools, by highlighting high-risk contexts even before the code is reviewed. We make our replication package publicly available to support further research and validation of our findings [29].
II Methodology
@@
Figure 1 shows an overview of our methodology. Using GitHub issues from SWE-bench, we collect developer patches and candidate patches generated by a standalone LLM (Llama 3.3) and three agentic frameworks (OpenHands, AutoCodeRover, HoneyComb). We analyze the prevalence and patterns of vulnerabilities for each of these code generation sources (RQ1, RQ2). In addition, we identify code, issue, and project-level factors that associated with the presence of these vulnerabilities (RQ3).
II-A Dataset and Task Setup
@@
We leverage the SWE-bench dataset [23], which consists of real-world GitHub issues paired with pull requests that resolve them. SWE-bench includes a train set of approximately 19,000 issue–PR pairs collected from 37 popular Python repositories and a test set of 2,294 issue–PR pairs from 12 repositories.
@@
For this study, we analyze two types of generated patches: those produced by a standalone LLM (Llama) and those generated by three top-performing agentic LLM frameworks (OpenHands, AutoCodeRover, HoneyComb).For the standalone model, we use Llama 3.3 Instruct (70B) to generate patches for both the train and test sets of SWE-bench, comprising approximately 21,000 instances in total. Since Llama is open-source and possible to run at scale, this generation step, although computationally heavy, is feasible. In contrast, for the agentic frameworks, we rely on publicly released patches that are available only for the test set. Due to the substantial computational and financial cost required to run these iterative frameworks across the train set, we restrict our framework-based analysis to the test split.
@@
Including both standalone LLM and agentic framework settings ensures the correct scope for addressing our research questions (RQ1–RQ2). These two settings also reflect distinct real-world usage scenarios. Standalone LLMs can better represent how developers interact with models, prompting them to get help with issue resolution and manually integrating the generated code. This makes the RQ1 analysis relevant for understanding the security implications of everyday developer usage. Agentic frameworks, in contrast, embody more autonomous approaches to issue resolution and reflect a growing trend toward automated, end-to-end program repair. Evaluating both settings enables a broader understanding of the security risks posed by current and emerging LLM-based workflows.
@@
Figure 1: Overview of Our Methodology for LLM-Based Code Generation and Security Evaluation.
II-B Patch Generation with LLM and Agentic Frameworks
II-B1 RQ1. Llama 3.3 Instruct (70B)
@@ II-B Patch Generation with LLM and Agentic Frameworks
To examine the security risks of patches produced by standalone LLMs, we generate candidate patches across the entire dataset (train+test) using the Llama 3.3 Instruct 70B model [24]. Recent benchmarks indicate that this model’s code generation capabilities are comparable to those of GPT-4o [30] and exceed the performance of Claude 3 Opus [31] and GPT-4 Turbo [32], making it a strong choice for our study [33, 34].
@@
Bug Localization: We use the oracle retrieval method, i.e., supplying the model with the corresponding issue description and the exact project files modified in the developer’s ground-truth fix. This ensures that the model receives the most relevant parts of the codebase that are needed to address the issue. We intentionally adopt the oracle retrieval method since it allows us to isolate and evaluate the security of the code generated by the LLM itself, without conflating generation quality with the efficacy of retrieval methods. Additionally, the default SWE-bench retrieval baseline, BM25, fails to retrieve any relevant files for approximately 40% of test instances [23]. The focus of our study is on assessing the security of generated code when an LLM is provided with the correct context.
@@
LLM Prompts: We adopt the prompt structure used in SWE-Bench, with slight modifications. Rather than generating diffs like SWE-bench, we explicitly instruct the model to produce complete files. While diff generation is popular for its lower token overhead and ease of use with patching tools [35], in standalone LLM settings, it often yields syntactically invalid or partially applicable patches, even after post-processing [23]. Our preliminary experiments confirmed this limitation. In contrast, producing full files ensures self-contained outputs without relying on diff syntax. It also reflects realistic usage, where developers expect standalone models to produce runnable code, not diffs. For this study, we prioritize reliability over minimal token usage.
II-B2 RQ2. Agentic LLM Frameworks
@@
To investigate whether agentic workflows introduce security risks in LLM-generated patches, we analyze solutions from three of the top-performing frameworks on the SWE-bench leaderboard: OpenHands+CodeAct v2.1 (Claude-3.5-Sonnet-20241022) [6], AutoCodeRover-v2.0 (Claude-3.5-Sonnet-20241022) [8], and Honeycomb [25]. These models resolved 29.38%, 24.89%, and 22.06% of the SWE-bench test set issues, respectively. We selected them based on their resolution rates at the time of our study and the availability of their generated patches on the SWE-bench official repositories [36]. Hereafter, we refer to these models as OH, ACR, and HC.
@@
II-C Developer-written Patch Collection
@@
To create a security baseline for comparison with LLM-generated code, we collect the original developer-written solutions for each issue from the corresponding PR using the GitHub API [37]. For each instance, we retrieve the final version of the files modified or created by the developer as part of the PR and merged into the codebase.
II-D Vulnerability Detection Using Majority Voting
@@
Manually analyzing the security of 20,000+ patches is infeasible. Therefore, we rely on three established OWASP-recommended [38] static analysis tools to automate the vulnerability detection process. The following tools were selected based on their widespread use [18, 39, 12], support for both rule-based and taint analysis, and their capacity to detect a wide range of real-world vulnerabilities [40]:
@@
CodeQL [28] is a static analysis tool developed by GitHub that is often used to identify security issues and code quality problems through heuristic and taint analysis techniques.
@@
Semgrep [26] is a static analysis tool that uses pattern-based rules and taint analysis to detect a wide range of vulnerabilities, e.g., injection risks and insecure API usage.
@@
Bandit[27] is a heuristic-based and security-focused static analysis tool designed for Python projects. Bandit inspects abstract syntax trees to detect common security issues.
A common concern with static analysis tools is their limited accuracy, particularly their tendency to produce false positives [41, 42, 43]. Therefore, to improve reliability and mitigate tool-specific false positives, we adopt a majority voting strategy: a vulnerability is accepted if at least two of the three tools detect it in the same file. Prior work has shown that static analysis tools often report distinct sets of vulnerabilities with limited overlap, and some studies have explored combining results via union or hybrid strategies to improve coverage [44, 45]. In contrast, our approach prioritizes precision over recall; by requiring agreement among tools, we aim to reduce false positives and improve the confidence in our findings. While this majority-vote strategy may miss some true positives, it offers a more conservative and reliable performance.
We focus on identifying only the new vulnerabilities introduced by LLM-generated code. By comparing original and modified files, we distinguish vulnerabilities introduced by the LLM from those already present. This allows us to isolate the security risks attributable to LLM-generated patches.
We apply the same static analysis tools and majority vote aggregation process to detect vulnerabilities across all sources, LLMs, agentic frameworks, and developer-written patches.
II-E Metrics for Empirical Analysis

### v3

# How Safe Are AI-Generated Patches? A Large-scale Study on Security Risks in LLM and Agentic Automated Program Repair on SWE-bench

@@ Abstract.
Large language models (LLMs) and their agentic frameworks are increasingly adopted to perform development tasks such as automated program repair (APR). While prior work has identified security risks in LLM-generated code, most have focused on synthetic, simplified, or isolated tasks that lack the complexity of real-world program repair. In this study, we present the first large-scale security analysis of LLM-generated patches using 20,000+ GitHub issues. We evaluate patches proposed by developers, a standalone LLM (Llama 3.3 Instruct-70B), and three top-performing agentic frameworks (OpenHands, AutoCodeRover, HoneyComb). Finally, we analyze a wide range of code, issue, and project-level factors to understand the conditions under which generating insecure patches is more likely.
Our findings reveal that Llama introduces many new vulnerabilities, exhibiting unique patterns not found in developers’ code. Agentic workflows also generate a number of vulnerabilities, particularly when given more autonomy. We find that vulnerabilities in LLM-generated patches are associated with distinctive code characteristics and are commonly observed in issues missing specific types of information. These results suggest that contextual factors play a critical role in the security of the generated patches and point toward the need for proactive risk assessment methods that account for both issue and code-level information.
LLMs, Agentic Frameworks, Security, Automated Program Repair
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacyccs: Security and privacy Software and application securityccs: Computing methodologies Natural language processing
1. Introduction
@@
LLMs are rapidly transforming software development, assisting with tasks like code completion, test generation, and bug fixing (li2023cctest; siddiq2024using; zhang2023well; alshahwan2024automated; wei2023copiloting). They are also increasingly integrated into agentic frameworks that enable autonomous planning and execution of complex tasks, such as automated program repair (APR), where LLM-based agents identify, diagnose, and patch software bugs directly in real-world projects (openhands; yang2024sweagent; zhang2024autocoderover). These systems have significantly outperformed traditional APR approaches and are now used to solve repairs across numerous real-world software projects, spanning thousands of issues from production-grade libraries and frameworks. This widespread adoption of LLMs in APR systems raises critical concerns about their potential security implications (perry2023users; yao2024survey; sajadi2025llms; siddiq2023generate). Studies show that 30-50% of code generated with LLM-based code assistants contains vulnerabilities (pearce2025asleep; fu2023security; wang2024aigeneratedcodereallysafe). When embedded within complex automated program repair frameworks, LLM-generated patches are often pushed directly into CI/CD pipelines, where a single insecure repair can rapidly propagate across dependent services and compromise entire software stacks.
While several studies have shown that LLMs are prone to generating insecure or low-quality code, most focus on simplified tasks such as coding exercises or identifying vulnerabilities in isolated code snippets (mohsin2024can; siddiq2024quality; pearce2025asleep; khoury2023secure). These scenarios lack the complexity and context of program repair in real projects, where understanding environment constraints or project-specific dependencies is critical. Moreover, such scenarios use LLMs as generic code generators rather than APR systems that produce targeted edits for specific bugs. APR agents also operate in a different setting than ordinary code agents. Code AI agents generate code from a short user prompt, while APR agents take a bug report, inspect the current repository state, and then propose candidate patches to fix a specific bug (openhands; zhang2024autocoderover). Hence, assessing their security requires evaluation frameworks that capture the full end-to-end repair workflow rather than isolated snippet-based analyses.
Although recent benchmarks like SWE-bench (jimenez2024swebench) have enabled the assessment of LLMs on real-world software engineering tasks, security-specific analyses in these contexts remain limited, and we still lack a systematic understanding of when APR-style LLM and agentic frameworks introduce new vulnerabilities, or how those vulnerabilities relate to properties of the patches and the surrounding development context. Recent studies have demonstrated that adversarial bug reports can mislead APR agents into producing insecure yet functionally correct patches (przymus2025adversarialbugreportssecurity; chen2025redteamingprogramrepair). These efforts evaluate security risks only assuming adversarial intent and on relatively small datasets. No prior work has performed a large-scale and security-focused empirical analysis of patches generated by LLM and agentic APR systems, especially assessing if vulnerabilities in patches arise even without adversarial intent.
Using SWE-bench, we analyze patches generated by Llama 3.3 and three state-of-the-art agentic APR frameworks, as well as the developer-written fixes for the same issues. We first quantify how often these systems introduce new vulnerabilities and which CWE types they tend to produce. We then examine how code-level, issue-level, and project-level factors influence the security of the proposed solutions, identifying conditions that result in vulnerability-inducing AI-generated patches. Our overarching goal with this work is to identify when and where APR systems are most likely to generate vulnerable patches, informing guidelines for when to trust their output and when to require additional checks. Overall, these analyses provide, to our knowledge, the first comprehensive evidence of the conditions and the frequency with which APR systems introduce security vulnerabilities when resolving real GitHub issues. Specifically, we ask the following research questions:
@@
  • RQ1: How secure are the patches produced by Llama 3.3 on real-world issue resolution tasks? How does their security compare to developer-written patches?
@@
Our findings highlight the security risks of LLM-generated code at scale. When evaluated on 20,000+ GitHub issues, the standalone Llama introduced 135 new vulnerabilities i.e., nearly 11x higher than the 12 new vulnerabilities introduced by developers. On a smaller test set of 2,200+ issues, the agentic frameworks also produced vulnerabilities, with the highest counts observed in the framework that gave the LLM greatest autonomy. Notably, some of the most recurring vulnerabilities introduced by LLMs (CWE-95, CWE-502) were distinct from those written by developers (e.g., CWE-732, CWE-377), suggesting that LLMs are not only replicating developers’ errors but also exhibiting new behavioral patterns. We found that vulnerable outputs are more likely when LLM patches involve a higher number of files and unique file type modifications. Further, most standalone LLM-generated vulnerabilities were observed in instances involving bug-related issues that do not contain code examples, information about the expected behavior of the code, or steps to reproduce the issue. These findings suggest that the risk of generating vulnerabilities is not only a function of the code produced but also of the context in which the LLM operates. As such, our insights can help inform proactive risk assessments of LLM and agent contributions, complementing the existing vulnerability detection tools, by highlighting high-risk repair tasks even before the code is generated or reviewed.
2. Methodology
@@
Figure 1 shows an overview of our methodology. Using GitHub issues from SWE-bench, we collect developer patches and candidate patches generated by a standalone LLM (Llama 3.3 (meta2025llama3)) and three agentic frameworks (OpenHands (openhands), AutoCodeRover (zhang2024autocoderover), HoneyComb (honeycomb)). We analyze the prevalence and patterns of vulnerabilities for each of these code generation sources (RQ1, RQ2). In addition, we identify code, issue, and project-level factors that associated with the presence of these vulnerabilities (RQ3). Our data and research artifacts publicly accessible through our replication package to support further research and validation of our findings (llm_security2025).
2.1. Dataset and Task Setup
@@
We leverage the SWE-bench dataset (jimenez2024swebench), which consists of real-world GitHub issues paired with pull requests that resolve them. SWE-bench includes a train set of approximately 19,000 issue–PR pairs collected from 37 popular Python repositories and a test set of 2,294 issue–PR pairs from 12 repositories.
@@
For the standalone model, we use Llama 3.3 Instruct (70B) to generate patches for both the train and test sets of SWE-bench, comprising approximately 21,000 instances in total. Since Llama is open-source and possible to run at scale, this generation step, although computationally heavy, is feasible. In contrast, for the agentic frameworks, we rely on publicly released patches that are available only for the test set. Due to the substantial computational and financial cost required to run these iterative frameworks across the train set, we restrict our framework-based analysis to the test split.
@@
Including both standalone LLM and agentic framework settings ensures the correct scope for addressing our research questions (RQ1–RQ2). These two settings also reflect distinct real-world usage scenarios. Standalone LLMs can better represent how developers interact with models, prompting them to get help with issue resolution and manually integrating the generated code. This makes the RQ1 analysis relevant for understanding the security implications of everyday developer usage. Agentic frameworks, in contrast, embody more autonomous approaches to issue resolution and reflect a growing trend toward automated, end-to-end program repair. Evaluating both settings enables a broader understanding of the security risks posed by current and emerging LLM-based workflows. Note that although any LLM-generated code can be further revised by developers, in line with SWE-bench and related work, our evaluation isolates the LLM’s own contributions. This allows us to assess the baseline security performance of the models and agents “out of the box,” without confounding effects from human review or refinement.
@@
Figure 1. Overview of Our Methodology for LLM-Based Code Generation and Security Evaluation.
2.2. Patch Generation with LLM and Agentic Frameworks
2.2.1. RQ1. Llama 3.3 Instruct (70B)
@@ 2.2. Patch Generation with LLM and Agentic Frameworks
To examine the security risks of patches produced by a standalone LLM, we generate candidate patches across the entire dataset (train+test) using the Llama 3.3 Instruct 70B model (meta2025llama3). Recent benchmarks indicate that this model’s code generation capabilities are comparable to those of GPT-4o (openai2025gpt4o) and exceed the performance of Claude 3 Opus (anthropic2024claude3) and GPT-4 Turbo (openai2023gpt4turbo), making it a strong choice for our study (lmarena2024leaderboard; chiang2024chatbot). We select this specific model over the older Code Llama family because published benchmark results show that Llama 3.3 outperforms the Code Llama variants. Further, due to the computational cost of generating patches for over 20,000 issues, evaluating additional 70B-scale models is not feasible.
@@
Bug Localization: We use the oracle retrieval method, i.e., supplying the model with the corresponding issue description and the exact project files modified in the developer’s ground-truth fix. This ensures that the model receives the most relevant parts of the codebase that are needed to address the issue. We intentionally adopt the oracle retrieval method since it allows us to isolate and evaluate the security of the code generated by the LLM itself, without conflating generation quality with the efficacy of retrieval methods. Additionally, the default SWE-bench retrieval baseline, BM25, fails to retrieve any relevant files for approximately 40% of test instances (jimenez2024swebench). The focus of our study is on assessing the security of generated code when an LLM is provided with the correct context.
@@
LLM Prompts: We adopt the prompt structure used in SWE-bench, with slight modifications. Rather than generating diffs like SWE-bench, we explicitly instruct the model to produce complete files. While diff generation is popular for its lower token overhead and ease of use with patching tools (wang2025solved), in standalone LLM settings, it often yields syntactically invalid or partially applicable patches, even after post-processing (jimenez2024swebench). Our preliminary experiments confirmed this limitation. In contrast, producing full files ensures self-contained outputs without relying on diff syntax. It also reflects realistic usage, where developers expect standalone models to produce runnable code, not diffs. For this study, we prioritize reliability over minimal token usage.
2.2.2. RQ2. Agentic LLM Frameworks
@@
To investigate whether agentic workflows introduce security risks in LLM-generated patches, we analyze solutions from three of the top-performing frameworks on the SWE-bench leaderboard: OpenHands+CodeAct v2.1 (Claude-3.5-Sonnet-20241022) (openhands), AutoCodeRover-v2.0 (Claude-3.5-Sonnet-20241022) (zhang2024autocoderover), and HoneyComb (honeycomb). These framework-model combinations resolved 29.38%, 24.89%, and 22.06% of the SWE-bench test set issues, respectively. As the top publicly available entries on the leaderboard at the time of our study, they represent the strongest functional performance that agentic workflows could achieve under this benchmark (swebench2023experiments). Hereafter, we refer to these models as OH, ACR, and HC.
@@
It is important to emphasize that our RQ2 evaluation is not a direct comparison between agentic frameworks and Llama. OH, ACR, and HC are evaluated using the test split, whereas the Llama and developer patches are analyzed on the combined train+test set. Our research questions are intentionally scoped to characterize behavior within each setting. For Llama, we measure how often its patches introduce vulnerabilities compared to developers on the same issues; for the agentic frameworks, we ask whether they also introduce vulnerabilities on the subset of issues they attempt to resolve, and identify the kinds of vulnerabilities they introduce through their patches. We therefore restrict all quantitative conclusions to the datasets on which each method is actually evaluated and avoid any claims of statistical superiority or inferiority between Llama and the frameworks.
2.3. Developer-written Patch Collection
@@
To create a security baseline for comparison with LLM-generated code, we collect the original developer-written solutions for each issue from the corresponding PR using the GitHub API (github_api). For each instance, we retrieve the final version of the files modified or created by the developer as part of the PR and merged into the codebase.
2.4. Vulnerability Detection
@@
Manually analyzing the security of 20,000+ patches across multiple sources is infeasible. Therefore, in line with prior studies (siddiq2024sallm; hamer2024just; siddiq2024quality; della2025prompt), we utilize established OWASP-recommended (owasp_source_code_analysis) static analysis tools to automate the vulnerability detection process. The following tools were selected based on their widespread use (siddiq2024quality; kree2024using; siddiq2023generate), support for both rule-based and taint analysis, and their capacity to detect a wide range of real-world vulnerabilities (li2025iris):
@@
CodeQL (github_codeql) is a static analysis tool developed by GitHub that is often used to identify security issues and code quality problems through heuristic and taint analysis techniques.
@@
Semgrep (semgrep) is a static analysis tool that uses pattern-based rules and taint analysis to detect a wide range of vulnerabilities, e.g., injection risks and insecure API usage.
@@
Bandit(bandit) is a heuristic-based and security-focused static analysis tool designed for Python projects. Bandit inspects abstract syntax trees to detect common security issues.
When analyzing patches, we run each static analysis tool on both the original (pre-patch) and modified (post-patch) files. By comparing the two, we classify findings as either persisting vulnerabilities (present in both the original and modified code) or new vulnerabilities (introduced only after the patch). Only the latter are counted in our results, as they directly capture the security risks added as a result of the modifications.
A common concern with static analysis tools is their accuracy, particularly their tendency to produce false positives (medeiros2014automatic; guo2023mitigating; dimastrogiovanni2016towards; nadeem2012high). To overcome any shortcomings in the recall and precision of static analysis tools, we create a multi-step process as follows:
Step 1: To maximize coverage (recall), we combined the results of three static analysis tools, in line with prior studies (shahid2025llm), since it has been shown that different tools have complementary strengths and that combining them provides a more comprehensive assessment of code quality (nunes2017combining; trautsch2023automated; charoenwet2024empirical; nunes2015phpsafe). The findings of the static analysis tools are aggregated using their reported vulnerability types (CWEs). Using multiple tools in this way improves coverage and recall in two ways: (A) it broadens the set of CWE types we can observe, since any single tool is limited to the classes of vulnerabilities it supports, and (B) even when all three tools nominally cover the same CWE, any one of them can miss a concrete instance that another flags due to differences in internal analyses. All three tools perform pattern-based static analysis, but CodeQL and Semgrep additionally support taint analysis (relying on out-of-the-box taint specifications for popular APIs), and CodeQL can perform partial constraint reasoning. These complementary analyses mean that aggregating results across tools yields the highest number of true positive findings overall, thereby maximizing recall. Here, the file-level focus of the analyses not only distinguishes between new and pre-existing issues but also helps reconcile the fact that static analysis tools can report inconsistent information (e.g., severity levels, line ranges, or explanations) for the same underlying vulnerabilities (lenarduzzi2023critical).
Step 2: To increase confidence (precision), we adopt a conservative majority voting strategy: a generated or developer-written patch is treated as a security risk only if at least two of the three tools independently identify new vulnerabilities in the same file(s) it modifies or creates. Moving beyond aggregation, the majority vote dismisses findings reported by only a single tool as likely false positives. Prior work has shown that a higher concentration of static analysis warnings correlates with an increased likelihood of risky code, and that considering such concentrations can improve precision in practice (trautsch2023automated; charoenwet2024empirical). Additionally, merging results from multiple static analysis tools has been shown to significantly reduce the likelihood of incorrect detections (dyremark2025false; meng2008approach; algaith2018finding). Accordingly, independent agreement across tools both strengthens the case that a patch should be treated with caution and reduces false positives, thereby improving precision while retaining the broader coverage benefits of combining multiple tools. This conservative approach can provide a balanced and trustworthy basis for our large-scale comparative analysis across different patch sources, including developer patches, standalone LLMs, and agentic frameworks.
Step 3: To ensure validity of all instances and to guarantee that only valid security-relevant issues enter our analysis, we performed a manual inspection in which the first author inspected the patched file, and judged whether the tool-reported issue constituted a plausible security threat rather than a benign code smell. Throughout the manual inspection, whenever a finding occurred inside a function, we conservatively assumed attacker-controlled input could reach the function. The instance was retained only if no adequate input validation or sanitization was evident within the function, along the path to the relevant vulnerable lines (e.g., whitelists, length/range checks, special-character checks, or calls to dedicated sanitizers). We excluded cases where inputs were hard-coded to safe values in one-off scripts or otherwise not reusable as library functionality. We also evaluated the security of instances in light of intent. For example, instances labeled for use of SHA-1 were removed when hashing served non-cryptographic purposes (e.g., content addressing). Conversely, we retained findings when untrusted parameters could influence critical operations, such as command execution (shell=True with function arguments), deserialization, path construction, or SQL string creation, and no adequate sanitization was evident.
This multi-step process addresses known limitations of fully automated static analysis and ensures the quality of our findings prior to our quantitative analyses.
2.5. Metrics for Empirical Analysis